# Lennart Berg and Torben Wiedemann

### v1

@@ 8 The blueprint computation for $H_{3}$-graded groups
For all $x,y\in\mathcal{S}$, we have
$[\theta_{-\delta}(x),\theta_{\gamma}(y)]$ $=\theta_{-\varepsilon}\big{(}h_{1}(x^{*},-y)^{*}\big{)}\theta_{\alpha}\big{(}-h_{2}(x^{*},-y)^{*}\big{)}\theta_{\beta}\big{(}-h_{1}(y,x^{*})\big{)}.$
@@
For all $x,y\in\mathcal{S}$, we know from 8.10 that
$[\theta_{\alpha}(x),\theta_{\varepsilon}(y)]=\theta_{\beta}\big{(}h_{1}(x,y)\big{)}\theta_{\gamma}\big{(}h_{2}(x,y)\big{)}\theta_{\delta}\big{(}-h_{1}(-y,x)\big{)}.$
Using the values of the parity map $\eta$ from Figure 5 and (6.20), we can conjugate this equation by $w_{\alpha}w_{\varepsilon}$. This yields
$[\theta_{-\delta}(x^{*}),\theta_{\gamma}(-y)]=\theta_{-\varepsilon}\big{(}h_{1}(x,y)^{*}\big{)}\theta_{\alpha}\big{(}-h_{2}(x,y)^{*}\big{)}\theta_{\beta}\big{(}-h_{1}(-y,x)\big{)}$
for all $x,y\in\mathcal{S}$. The assertion follows. ∎
@@
Let $(\bar{\alpha},\bar{\beta},\bar{\gamma},\bar{\delta},\bar{\varepsilon})$ be an arbitrary $H_{2}$-quintuple. Then $\theta_{\bar{\alpha}}(\mathcal{S}_{2})$ and $\theta_{\bar{\delta}}(\mathcal{S}_{2})$ commute.
@@
We have $g(\mathcal{S}_{2},\mathcal{S}_{2})=\{0\}$ by 8.29 ii. Thus by 8.10 v, the assertion holds for $(\bar{\alpha},\bar{\beta},\bar{\gamma},\bar{\delta},\bar{\varepsilon})=(\alpha,\beta,\gamma,\delta,\varepsilon)$. The general assertion follows by conjugation with Weyl elements. ∎
@@
For all $\alpha\in H_{3}$, we have $[\theta_{\alpha}(\mathcal{S}_{1}),\theta_{-\alpha}(\mathcal{S}_{2})]=\{1\}$.
@@
Let $x,y\in\mathcal{S}_{1}$. By 8.35, we have $ABC=1_{G}$ where
$A$ $=\big{[}[\theta_{-\delta}([0,1]),\theta_{\gamma}([0,x])],\theta_{-\alpha}([0,y])^{\theta_{-\delta}([0,1])}],$
$B$ $=\big{[}[\theta_{-\alpha}([0,y]),\theta_{-\delta}([0,1])],\theta_{\gamma}([0,x])^{\theta_{-\alpha}([0,y])}],$
$C$ $=\big{[}[\theta_{\gamma}([0,x]),\theta_{-\alpha}([0,y])],\theta_{-\delta}([0,1])^{\theta_{\gamma}([0,x])}].$
It follows from 8.34 that $B=1=C$ and that $\theta_{-\alpha}([0,y])^{\theta_{-\delta}([0,1])}=\theta_{-\alpha}([0,y])$. Further, since $h_{1}(\mathcal{S}_{2},\mathcal{S}_{2})=\{0\}$ by 8.29 iii, we infer from 8.33 that
$[\theta_{-\delta}([0,1]),\theta_{\gamma}([0,x])]$ $=\theta_{\alpha}\big{(}-h_{2}([0,1],[0,-x])^{*}\big{)}=\theta_{\alpha}([x,0]).$
@@
$1_{G}$ $=ABC=A=[\theta_{\alpha}([x,0]),\theta_{-\alpha}([0,y])].$
@@
We now define the $D_{6}$-grading of $G$ whose folding is $(U_{\alpha})_{\alpha\in H_{3}}$.
Notation 8.37.
For the rest of this section, we use the notation of 4.5 and 4.6. In particular, we write $\Delta_{D}=(\delta_{1},\ldots,\delta_{6})$ for the standard root base of $D_{6}$, $\Delta_{H}=\Delta=(\rho_{1},\rho_{2},\rho_{3})$ for the (corresponding) standard root base of $H_{3}$ and $\phi\circ\pi$ for the chain of maps $D_{6}\rightarrow GH_{3}\rightarrow H_{3}$. For all $\alpha\in H_{3}$, we define $V_{\kappa^{\text{sh}}(\alpha)}\coloneqq\theta_{\alpha}(\mathcal{S}_{1})$ and $V_{\kappa^{\text{lo}}(\alpha)}\coloneqq\theta_{\alpha}(\mathcal{S}_{2})$ where $\kappa^{\text{sh}}(\alpha),\kappa^{\text{lo}}(\alpha)\in D_{6}$ are as in 4.6. In this way, we have defined a root group $V_{\beta}$ for all $\beta\in D_{6}$.
@@
Let $(i,j)\in\{(1,6),(2,4),(3,5)\}$ and put $\rho\coloneqq\phi\circ\pi(\delta_{i})=\phi\circ\pi(\delta_{j})\in\Delta\subset H_{3}$. Then for all $\xi\in D_{6}$, we have $V_{\xi}^{w_{\rho}}=V_{\xi^{\sigma(\delta_{i})\sigma(\delta_{j})}}$.
@@
Let $\zeta\in H_{3}$ and put $\xi\coloneqq\kappa^{\text{sh}}(\zeta)\in D_{6}$. Then
$V_{\xi}^{w_{\rho}}$ $=\theta_{\zeta}(\mathcal{S}_{1})^{w_{\rho}}=\theta_{\zeta^{\sigma(\rho)}}(\eta_{\zeta,\rho}.\mathcal{S}_{1})=\theta_{\zeta^{\sigma(\rho)}}(\mathcal{S}_{1})=V_{\kappa^{\text{sh}}(\zeta^{\sigma(\rho)})}$
$=V_{\kappa^{\text{sh}}(\zeta)^{\sigma(\delta_{i})\sigma(\delta_{j})}}=V_{\xi^{\sigma(\delta_{i})\sigma(\delta_{j})}}.$
by 4.13. The assertion for $\xi\coloneqq\kappa^{\text{lo}}(\zeta)$ can be proven similarly. ∎
@@
Let $\xi_{1},\xi_{2}\in D_{6}$ and put $\zeta_{1}\coloneqq\phi\circ\pi(\xi_{1}),\zeta_{2}\coloneqq\phi\circ\pi(\xi_{2})\in H_{3}$. Assume that $\zeta_{1},\zeta_{2}\in\{\alpha,\beta,\gamma,\delta,\varepsilon\}$ or $\zeta_{1},\zeta_{2}\in\{\rho_{1},\rho_{1}+\rho_{2},\rho_{2}\}$ or $\zeta_{1},\zeta_{2}\in\{\rho_{1},\rho_{3}\}$. Then $[V_{\xi_{1}},V_{\xi_{2}}]\leq V_{\xi_{1}+\xi_{2}}$ if $\xi_{1}+\xi_{2}$ is a root (in $D_{6}$) and $[V_{\xi_{1}},V_{\xi_{2}}]=\{1\}$ otherwise.
@@
The $D_{6}$-root groups $V_{\xi_{1}},V_{\xi_{2}}$ are contained in the $H_{3}$-root groups $U_{\zeta_{1}}$, $U_{\zeta_{2}}$, respectively. For the latter, we have a commutator formula $[U_{\zeta_{1}},U_{\zeta_{2}}]\subseteq U_{\mathord{]\zeta_{1},\zeta_{2}[}}$. By 8.30, the commutator maps $\psi_{\zeta_{1},\zeta_{2}}^{\rho}$ in this commutator formula are of the same form as in the folding $(\operatorname{Ch}(D_{6},\mathbb{C}),(\tilde{U}_{\rho}^{D})_{\rho\in D_{6}})$. The assertion follows. ∎
@@
The family $(V_{\xi})_{\xi\in D_{6}}$ is a $D_{6}$-grading of $G$, and its folding with respect to the diagram in Figure 1(b) is precisely $(U_{\zeta})_{\zeta\in H_{3}}$. In particular, every $H_{3}$-graded group is the folding of a $D_{6}$-graded group.
@@
We begin by verifying the $D_{6}$-commutator relations. Let $\xi_{1},\xi_{2}\in D_{6}$ be non-proportional and put $\zeta_{1}\coloneqq\pi(\xi_{1})$, $\zeta_{2}\coloneqq\pi(\xi_{2})\in GH_{3}$. If $\zeta_{1}=e\zeta_{2}$ for some $e\in\{\pm 1\}$, then $\xi_{1}=e\xi_{2}$ because $\pi\colon D_{6}\rightarrow GH_{3}$ is bijective and linear, which contradicts the assumption that $\xi_{1},\xi_{2}$ are non-proportional. If $\zeta_{1},\zeta_{2}$ are proportional but not of the same length, then $[V_{\xi_{1}},V_{\xi_{2}}]=\{1_{G}\}$ by 8.36. Hence we can assume that $\zeta_{1},\zeta_{2}$ are non-proportional.
Choose a root base $\Delta_{H}^{\prime}$ of the rank-2 subsystem of $H_{3}$ spanned by $\bar{\zeta}_{1}\coloneqq\phi(\zeta_{1})$ and $\bar{\zeta}_{2}\coloneqq\phi(\zeta_{2})$ such that $\bar{\zeta}_{1},\bar{\zeta}_{2}$ are $\Delta_{H}^{\prime}$-positive and put $r_{1}\coloneqq\sigma(\delta_{1})\sigma(\delta_{6})$, $r_{2}\coloneqq\sigma(\delta_{2})\sigma(\delta_{4})$, $r_{3}\coloneqq\sigma(\delta_{3})\sigma(\delta_{5})$. Recall from 4.5 and 4.13 that $r_{1},r_{2},r_{3}$ are the simple reflections in $\operatorname{Weyl}(H_{3})$ corresponding to $\Delta_{H}=(\rho_{1},\rho_{2},\rho_{3})$ and that they also act on $D_{6}$. The transitivity properties of the Weyl group yield $i_{1},\ldots,i_{k}\in\{1,2,3\}$ such that $\Delta_{H}$ contains $(\Delta_{H}^{\prime})^{r}$ where $r\coloneqq r_{i_{1}}\cdots r_{i_{k}}$. For all $j\in\{1,2\}$, we further put $\bar{\zeta}_{j}^{\prime}\coloneqq\bar{\zeta}_{j}^{r}$ and $\xi_{j}^{\prime}\coloneqq\xi_{j}^{r}$, so that
$\phi\pi(\xi_{j}^{\prime})=\phi(\pi(\xi_{j})^{r})=\phi(\zeta_{j}^{r})=\phi(\zeta_{j})^{r}=\bar{\zeta}_{j}^{\prime}$
by 4.13. Then $\bar{\zeta}_{1}^{\prime}$ and $\bar{\zeta}_{2}^{\prime}$ are $\Delta_{H}$-positive, so it follows from 8.39 that $V_{\xi_{1}^{\prime}}$ and $V_{\xi_{2}^{\prime}}$ satisfy the desired commutator relations. By conjugation with $(w_{\rho_{i_{1}}}\cdots w_{\rho_{i_{k}}})^{-1}$ and an application of 8.38, it follows that $V_{\xi_{1}}$ and $V_{\xi_{2}}$ satisfy the desired commutator relations as well. Hence $(V_{\xi})_{\xi\in D_{6}}$ satisfies $D_{6}$-commutator relations.

### v2

@@ 10 All $H_{\ell}$-gradings arise from foldings
Denote by $(\alpha,\beta,\gamma,\delta,\varepsilon)$ the $H_{2}$-quintuple associated to $(\rho_{2},\rho_{3})$. Then for all $x,y\in\mathcal{R}\times\mathcal{R}$, we have
$[\theta_{-\delta}^{H}(x),\theta_{\gamma}^{H}(y)]$ $=\theta_{-\varepsilon}^{H}\big{(}h_{1}(x^{*},-y)^{*}\big{)}\theta_{\alpha}^{H}\big{(}-h_{2}(x^{*},-y)^{*}\big{)}\theta_{\beta}^{H}\big{(}-h_{1}(y,x^{*})\big{)}.$
@@
For all $x,y\in\mathcal{R}\times\mathcal{R}$, we know from 9.10 that
$[\theta_{\alpha}^{H}(x),\theta_{\varepsilon}^{H}(y)]=\theta_{\beta}^{H}\big{(}h_{1}(x,y)\big{)}\theta_{\gamma}^{H}\big{(}h_{2}(x,y)\big{)}\theta_{\delta}^{H}\big{(}-h_{1}(-y,x)\big{)}.$
Using the values of the standard parity map $\eta$ and (6.25), we can conjugate this equation by $w_{\alpha}w_{\varepsilon}$. This yields
$[\theta_{-\delta}^{H}(x^{*}),\theta_{\gamma}^{H}(-y)]=\theta_{-\varepsilon}^{H}\big{(}h_{1}(x,y)^{*}\big{)}\theta_{\alpha}^{H}\big{(}-h_{2}(x,y)^{*}\big{)}\theta_{\beta}^{H}\big{(}-h_{1}(-y,x)\big{)}$
for all $x,y\in\mathcal{R}\times\mathcal{R}$. The assertion follows. ∎
@@
Let $(\bar{\alpha},\bar{\beta},\bar{\gamma},\bar{\delta},\bar{\varepsilon})$ be any $H_{2}$-quintuple in $H_{\ell}$. Then $\theta_{\bar{\alpha}}^{H}(\mathcal{R}_{2})$ and $\theta_{\bar{\delta}}^{H}(\mathcal{R}_{2})$ commute.
@@
We have $g(\mathcal{R}_{2},\mathcal{R}_{2})=\{0\}$ by 9.29 ii. Thus by 9.10 v, $\psi_{\alpha,\delta}^{\beta}(\mathcal{R}_{2},\mathcal{R}_{2})=\{0\}=\psi_{\alpha,\delta}^{\gamma}(\mathcal{R}_{2},\mathcal{R}_{2})$ where $(\alpha,\beta,\gamma,\delta,\varepsilon)$ denotes the $H_{2}$-quintuple associated to $(\rho_{2},\rho_{3})$. Hence the assertion holds for $(\bar{\alpha},\bar{\beta},\bar{\gamma},\bar{\delta},\bar{\varepsilon})=(\alpha,\beta,\gamma,\delta,\varepsilon)$. The general assertion follows by conjugation with suitable Weyl elements. ∎
@@
For all $\zeta\in H_{\ell}$, we have $[\theta_{\zeta}^{H}(\mathcal{R}_{1}),\theta_{-\zeta}^{H}(\mathcal{R}_{2})]=\{1\}$.
@@
Let $x,y\in\mathcal{R}$ and denote by $(\alpha,\beta,\gamma,\delta,\varepsilon)$ the $H_{2}$-quintuple associated to $(\rho_{2},\rho_{3})$. By conjugating with suitable Weyl elements, we may assume that $\zeta=\alpha$. By 10.4, we have $ABC=1_{G}$ where
$A$ $=\big{[}[\theta_{-\delta}^{H}([0,1]),\theta_{\gamma}^{H}([0,x])],\theta_{-\alpha}^{H}([0,y])^{\theta_{-\delta}^{H}([0,1])}],$
$B$ $=\big{[}[\theta_{-\alpha}^{H}([0,y]),\theta_{-\delta}^{H}([0,1])],\theta_{\gamma}^{H}([0,x])^{\theta_{-\alpha}^{H}([0,y])}],$
$C$ $=\big{[}[\theta_{\gamma}^{H}([0,x]),\theta_{-\alpha}^{H}([0,y])],\theta_{-\delta}^{H}([0,1])^{\theta_{\gamma}^{H}([0,x])}].$
It follows from 10.3 that $B=1=C$ and that $\theta_{-\alpha}^{H}([0,y])^{\theta_{-\delta}^{H}([0,1])}=\theta_{-\alpha}^{H}([0,y])$. Further, since $h_{1}(\mathcal{R}_{2},\mathcal{R}_{2})=\{0\}$ by 9.29 iii, we infer from 10.2 that
$[\theta_{-\delta}^{H}([0,1]),\theta_{\gamma}^{H}([0,x])]$ $=\theta_{\alpha}^{H}\big{(}-h_{2}([0,1],[0,-x])^{*}\big{)}=\theta_{\alpha}^{H}([x,0]).$
@@
$1_{G}$ $=ABC=A=[\theta_{\alpha}^{H}([x,0]),\theta_{-\alpha}^{H}([0,y])].$
@@
We now define the $X$-grading of $G$ whose folding is $(U_{\alpha}^{H})_{\alpha\in H_{\ell}}$.
Definition 10.6.
For all $\xi\in X$, we define a root homomorphism $\theta_{\xi}^{X}\colon(\mathcal{R},+)\rightarrow U_{\pi(\xi)}^{H}$ by $\theta_{\xi}^{X}(s)\coloneqq\theta_{\pi(\xi)}^{H}([s,0])$ if $\tilde{\pi}(\xi)$ is short in $GH_{\ell}$ and by $\theta_{\xi}^{X}(s)\coloneqq\theta_{\pi(\xi)}^{H}([0,s])$ if $\tilde{\pi}(\xi)$ is long in $GH_{\ell}$. Further, we put $U_{\xi}^{X}\coloneqq\theta_{\xi}^{X}(\mathcal{R})$.
@@
Let $(i,j)\in\{(1,6),(2,4),(3,5)\}$ (or possibly $(i,j)=(7,8)$ if $X=E_{8}$) and put $\rho\coloneqq\pi(\delta_{i})=\pi(\delta_{j})\in\Delta\subset H_{\ell}$. Then for all $\xi\in X$, we have $(U_{\xi}^{X})^{w_{\rho}}=U_{\xi^{\sigma(\delta_{i})\sigma(\delta_{j})}}^{X}$.
@@
Let $\xi\in X$. Put $k\coloneqq 1$ if $\tilde{\pi}(\xi)$ is short in $GH_{\ell}$ and $k\coloneqq 2$ otherwise, so that $U_{\xi}^{X}=\theta_{\pi(\xi)}^{H}(\mathcal{R}_{k})$. Then
$(U_{\xi}^{X})^{w_{\rho}}=\theta_{\pi(\xi)}^{H}(\mathcal{R}_{k})^{w_{\rho}}=\theta_{\pi(\xi)^{\sigma(\rho)}}^{H}(\eta_{\pi(\xi),\rho}.\mathcal{R}_{k})=\theta_{\pi(\xi)^{\sigma(\rho)}}^{H}(\mathcal{R}_{k}).$
Here $\pi(\xi)^{\sigma(\rho)}=\pi(\xi^{u(\sigma(\rho))})=\pi(\xi^{\sigma(\delta_{i})\sigma(\delta_{j})})$ by 4.8 e. Further, by 4.16, $\tilde{\pi}(\xi^{\sigma(\delta_{i})\sigma(\delta_{j})})$ is short in $GH_{\ell}$ if and only if $\tilde{\pi}(\xi)$ is, so that
$U_{\xi^{\sigma(\delta_{i})\sigma(\delta_{j})}}^{X}=\theta_{\pi(\xi^{\sigma(\delta_{i})\sigma(\delta_{j})})}^{H}(\mathcal{R}_{k}).$
Hence
$(U_{\xi}^{X})^{w_{\rho}}$ $=\theta_{\pi(\xi^{\sigma(\delta_{i})\sigma(\delta_{j})})}^{H}(\mathcal{R}_{k})=U_{\xi^{\sigma(\delta_{i})\sigma(\delta_{j})}}^{X},$
as desired. ∎
@@
Let $\xi_{1},\xi_{2}\in X$ and assume that $\pi(\xi_{1})\neq\pi(\xi_{2})$ and that $\pi(\xi_{1})$, $\pi(\xi_{2})$ both lie in $E_{2}^{+}(\Delta)$, the set defined in 2.8. If $\xi_{1}+\xi_{2}$ is a root (in $X$), then
$[\theta_{\xi_{1}}^{X}(r),\theta_{\xi_{2}}^{X}(s)]=\theta_{\xi_{1}+\xi_{2}}^{X}(c_{\xi_{1},\xi_{2}}rs)$
for all $r,s\in\mathcal{R}$ where $c_{\xi_{1},\xi_{2}}\in\{\pm 1\}$ is the same sign that appears in the Chevalley group $\operatorname{Ch}(X,\mathcal{R})$ in 4.31. If $\xi_{1}+\xi_{2}$ is not a root (in $X$), then $[U_{\xi_{1}}^{X},U_{\xi_{2}}^{X}]=\{1\}$.
@@
The $X$-root groups $U_{\xi_{1}}^{X},U_{\xi_{2}}^{X}$ are contained in the $H_{\ell}$-root groups $U_{\zeta_{1}}^{H}$, $U_{\zeta_{2}}^{H}$, respectively. For the latter, we have a commutator formula $[U_{\zeta_{1}}^{H},U_{\zeta_{2}}^{H}]\subseteq U_{\mathord{]\zeta_{1},\zeta_{2}[}}^{H}$. By 9.31, the commutator maps $\psi_{\zeta_{1},\zeta_{2}}^{\rho}$ in this commutator formula are of the same form as in $(\operatorname{Ch}(X,\mathcal{R}),(U_{\rho}^{D})_{\rho\in D_{6}})$. The assertion follows. ∎
Example 10.9.
We illustrate the argument in the proof of 10.8 by the example $\xi_{1}\coloneqq e_{2}-e_{3}$, $\xi_{2}\coloneqq e_{3}-e_{4}\in D_{6}$. Then $\pi(\xi_{1})=\rho_{2}$ and $\pi(\xi_{2})=\rho_{3}$ by Figure 3. Denote by $(\alpha,\beta,\gamma,\delta,\varepsilon)$ the $H_{2}$-quintuple associated to $(\rho_{2},\rho_{3})$. The commutator formulas in Figure 5, which hold by 9.33, yield that
$\big{[}\theta_{\xi_{1}}^{X}(a),\theta_{\xi_{2}}^{X}(b)\big{]}=\big{[}\theta_{\alpha}^{H}([a,0]),\theta_{\varepsilon}^{H}([0,b])\big{]}=\theta_{\delta}^{H}([ab,0])$
for all $a,b\in\mathcal{R}_{1}$. Since $\delta=[0,1,\tau]$, we read off from Figure 3 that the short root in $\phi^{-1}(\delta)$ is $\tilde{\pi}(e_{2}-e_{4})$. We conclude that $[U_{\xi_{1}}^{X},U_{\xi_{2}}^{X}]\leq U_{\xi_{1}+\xi_{2}}^{X}$.
Lemma 10.10.
The group $G$ has $X$-commutator relations with respect to the root groups $(U_{\xi}^{X})_{\xi\in X}$ from 10.6 and has the same commutation maps (with respect to $(\theta_{\xi}^{X})_{\xi\in X}$) as the Chevalley group $\operatorname{Ch}(X,\mathcal{R}_{1})$ (with respect to the root isomorphisms constructed in 4.26).
Proof.
Let $\xi_{1},\xi_{2}\in X$ be non-proportional and put $\zeta_{1}\coloneqq\tilde{\pi}(\xi_{1})$, $\zeta_{2}\coloneqq\tilde{\pi}(\xi_{2})\in GH_{3}$. If $\zeta_{1}=e\zeta_{2}$ for some $e\in\{\pm 1\}$, then $\xi_{1}=e\xi_{2}$ because $\tilde{\pi}\colon X\rightarrow GH_{\ell}$ is bijective and linear, which contradicts the assumption that $\xi_{1},\xi_{2}$ are non-proportional. If $\zeta_{1}$, $\zeta_{2}$ differ by a factor of $\tau$, then $U_{\xi_{1}}^{X}$ and $U_{\xi_{2}}^{X}$ are contained in the same (abelian) $H_{\ell}$-root group $U_{\pi(\xi_{1})}^{H}=U_{\pi(\xi_{2})}^{H}$, so that $[U_{\xi_{1}}^{X},U_{\xi_{2}}^{X}]=\{1_{G}\}$. If $\zeta_{1}$, $\zeta_{2}$ differ by a factor of $-\tau$, then $\pi(\xi_{1})=-\pi(\xi_{2})$ and $[U_{\xi_{1}}^{X},U_{\xi_{2}}^{X}]=\{1_{G}\}$ by 10.5. Hence we can assume that $\zeta_{1},\zeta_{2}$ are non-proportional.
If $\zeta_{1}$, $\zeta_{2}$ both lie in $E_{2}^{+}(\Delta)$, then the assertion holds by 10.8. In general, there exists (by 2.9) some $w\in\operatorname{Weyl}(H_{\ell})$ such that $\zeta_{1}^{w}$ and $\zeta_{2}^{w}$ both lie in $E_{2}^{+}(\Delta)$. Thus we can conjugate the commutator relation of the pair $(U_{\zeta_{1}^{w}}^{X},U_{\zeta_{2}^{w}}^{X})$ by a suitable sequence of the Weyl elements $(w_{\delta}^{\prime})_{\delta^{\prime}\in\Delta}$ to obtain the commutator relation of $(U_{\zeta_{1}}^{X},U_{\zeta_{2}}^{X})$. Since the parametrisations of $G$ and $\operatorname{Ch}(X,\mathcal{R}_{1})$ adhere to the same parity map (namely, the standard parity map for $H_{\ell}$), we infer that the commutator relation of $(U_{\zeta_{1}}^{X},U_{\zeta_{2}}^{X})$ has the desired form as well (cf. 7.2). ∎
@@
Let $\ell\in\{3,4\}$, let $X$ denote the root system in $\{D_{6},E_{8}\}$ of rank $2\ell$ and let $(U_{\zeta}^{H})_{\zeta\in H_{\ell}}$ be an $H_{\ell}$-grading of a group $G$. Then the family $(U_{\xi}^{X})_{\xi\in X}$ defined in 10.6 is an $X$-grading of $G$, and its folding (in the sense of 4.19) is precisely $(U_{\zeta}^{H})_{\zeta\in H_{\ell}}$.